HOW DO YOU SOLVE (TRUE OR FALSE) " IF YOU ROLL SIX SIDED DIE SIX TIMES, YOU WILL ROLL AN EVEN NUMBER AT LEAST ONCE.
Answer by oberobic(2304)

There are 3 even numbers and 3 odd numbers.
.
The chance you will roll and even number is 3/6 = 1/2 for each roll.
.
Each roll is an independent event, so you multiply the probabilities.
.
If you roll 6 times, then the change of rolling only odd numbers = (1/2)^6 = 1/64.
.
So there is a 1/64 chance that you will roll only odd numbers.
.
That means there is a 63/64 chance that you will roll at least 1 even number.
.
Those are good odds, and it is unlikely you would roll all odd numbers. But you cannot guarantee an even number will be rolled, no matter how many times to roll. Rare things do happen.
.
So the answer appears to be false.